# Decoding by Contrasting Knowledge: Enhancing LLMs’ Confidence on Edited Facts

### v1

@@ Abstract
The knowledge within large language models (LLMs) may become outdated quickly. While in-context editing (ICE) is currently the most effective method for knowledge editing (KE), it is constrained by the black-box modeling of LLMs and thus lacks interpretability. Our work aims to elucidate the superior performance of ICE on the KE by analyzing the impacts of in-context new knowledge on token-wise distributions. We observe that despite a significant boost in logits of the new knowledge, the performance of is still hindered by stubborn knowledge. Stubborn knowledge refers to as facts that have gained excessive confidence during pretraining, making it hard to edit effectively. To address this issue and further enhance the performance of ICE, we propose a novel approach termed Decoding by Contrasting Knowledge (DeCK). DeCK derives the distribution of the next token by contrasting the logits obtained from the newly edited knowledge guided by ICE with those from the unedited parametric knowledge. Our experiments consistently demonstrate that DeCK enhances the confidence of LLMs in edited facts. For instance, it improves the performance of LLaMA3-8B-instruct on MQuAKE by up to 219%, demonstrating its capability to strengthen ICE in the editing of stubborn knowledge. Our work paves the way to develop the both effective and accountable KE methods for LLMs.  [^1] The source code is available at: https://github.com/byronBBL/DeCK.
@@ 1 Introduction
Unlike the high computational resource burden incurred by retraining from scratch, knowledge editing (KE) (Sinitsin et al.,, 2020; De Cao et al.,, 2021; Zhu et al.,, 2020; Mitchell et al.,, 2022; Yao et al.,, 2023) have been proposed as an efficient means to update the knowledge of LLMs. They aim to edit knowledge by incrementally injecting or modifying facts.
As LLMs demonstrate increasingly powerful in-context learning capabilities, recent research (Madaan et al.,, 2022; Zhong et al.,, 2023; Zheng et al.,, 2023; Cohen et al.,, 2024; Wang et al.,, 2024; Mei et al.,, 2024) has delved into more easy and efficient methods for in-context editing (ICE), aiming to directly guide frozen LLMs in generating text with new knowledge through contextual prompts. Figure 1 (left) illustrates an example of successful editing using ICE. These ICE methods showcasing state-of-the-art performance without the need to alter internal model parameters, indicate the promising potential of modeling LLMs as black boxes for ICE guided by external contexts.
@@
However, as illustrated in Figure 1 (middle), there still exist deeply entrenched pieces of knowledge in LLMs that is difficult for ICE to modify, which we refer to as stubborn knowledge. We argue that LLMs, through extensive pre-training, have developed strong confidence in certain facts, making it difficult to alter these stubborn knowledge solely through external contextual prompts (Bi et al.,, 2024). Therefore, despite the fact that the sophisticated methods such as enhancing retrieval (Shi et al.,, 2024), checking conflict (Zhong et al.,, 2023), and guiding reasoning (Wang et al.,, 2024) can enhance the overall performance of ICE, relying on these external methods still cannot genuinely improve the foundational capability for editing individual stubborn knowledge.
In this work, we focus on enhancing state-of-the-art KE methods ICE to overcome the stubborn knowledge in LLMs. First, we observe the impact of the in-context new knowledge in ICE on LLMs from the perspective of LLMs’ token-level distributions. We find that incorporating this new knowledge significantly increases the predicted probability of generating edited facts during the decoding process. A deeper exploration of the failed cases reveals the reasons why stubborn knowledge is hard to edit. Despite the significant improvement in the logits of new knowledge achieved by ICE, there persists a small gap between new knowledge and parametric knowledge, where parametric knowledge refers to the original unedited knowledge in LLMs.
Building upon the insights gained from above observations, we introduce a new decoding technique called Decoding by Contrasting Knowledge (DeCK) to enhance LLMs’ confidence in edited facts for better editing of stubborn knowledge. DeCK consists of two components: (1) an editing enhancement module that improve attention to new knowledge, thus preventing it from being filtered out during contrastive decoding, and (2) a contrastive decoding strategy that compares the logical distributions after in-context editing with the original parametric logical distributions to predict the next token.
Overall, our contributions can be summarized by three points. First, as far as we know, we are the first to elucidate superior performance of ICE on the KE from a model interpretability perspective. Second, we find that stubborn knowledge significantly impacts the performance of ICE, and we propose DeCK to boost confidence in editing facts, enhancing ICE to overcome it. Third, extensive experiments on MQuAKE indicate that DeCK can effectively enhance the performance of ICE without altering the internal model or modifying external prompts. DeCK can be easily integrated into any ICE method as a decoding component to enhance editing capabilities. Our work paves the way to develop the both effective and accountable KE methods for LLMs.
@@ 3 Deep Insights into ICE through Decoding Perspectives
With $\phi(\cdot)$ replacing the affine layer to predict the probability of the next token over the vocabulary set $\mathcal{V}$, we can obtain a simpler representation of Equation (1). Given a sequence of tokens $\mathcal{X}_{E}=\{x_{1}^{(E)},x_{2}^{(E)},...,x_{m-1}^{(E)}\}$, which includes guidance of an editing prompt, such as "Assume Ellie Kemper is a citizen of Croatia" in Figure 1, we compute the probability of next token $x^{(E)}_{m}$ with editing guidance as follows:
@@
We can also represent the parametric probability distribution $\text{I\kern-1.49994ptP}^{B}(x^{(B)}_{n}|x^{(B)}_{<n})$ by considering only the token sequence $\mathcal{X}_{B}$ of length $n$ containing the original question prompt without any editing content. The distribution $\text{I\kern-1.49994ptP}^{E}(x^{(E)}_{m}|x^{(E)}_{<m})$ also reflects the feedback from the introduction of external knowledge, while $\text{I\kern-1.49994ptP}^{B}(x^{(B)}_{n}|x^{(B)}_{<n})$ solely represents the response of LLMs based on their parametric knowledge to the question.
@@ 3.1 How ICE Effectively Edits Knowledge in LLMs?
Although the ICE methods on LLMs (Zhong et al.,, 2023; Cohen et al.,, 2024; Wang et al.,, 2024) have demonstrated promising performance, they all rely on black-box modeling of LLMs for editing, and the internal mechanisms behind their effectiveness remain unclear. In this subsection, we delve into the intrinsic reasons behind the superior performance of ICE on KE.
We designed dedicated experiments to capture the logits output of knowledge that would be influenced by the edit. A striking observation in Figure 2 is that the introduction of new knowledge through ICE leads to a significant rightward shift in the probability distribution of the new knowledge, while the logits for parametric knowledge remain largely unchanged or decrease to some extent. This suggests that ICE significantly enhances the logits of new knowledge while having minimal impact on parametric knowledge. In addition, the number of top-ranked positions for new knowledge significantly increases after ICE, with the majority surpassing that of parametric knowledge. This indicates that the in-context new knowledge can improve the confidence of LLMs in editing facts, thereby prompting responses with the edited answers.
@@ 3.2 Inherent Challenges of Stubborn Knowledge
While ICE has significantly boosted the confidence of large models in new knowledge, we find that there are still instances where certain new knowledge ranks prominently but not as the top-1, as illustrated in Figure 3. We term this phenomenon "stubborn knowledge", which refers to cases where editing fails due to either an excessive confidence in existing parametric knowledge or insufficient confidence in new knowledge. The edit cases in Figure 3 deeply reveals the failed pattern for ICE in addressing stubborn knowledge, which happens when there is still an extremely small gap compared to the parametric knowledge after editing, despite the significant increase in new knowledge logits induced by the editing prompt. Taking the last case as an example, after editing, the new knowledge "English" lags behind the parametric knowledge "French" by only 0.516 in terms of logical distribution, illustrating how a minor gap leads to editing failure. This indicates the intrinsic reasons for the failure of black-box ICE methods to edit stubborn knowledge in LLMs in most cases.
@@ 4 DeCK: Enhancing LLMs’ Confidence on Edited Facts
Inspired by the observations in Section 3, we design our novel decoding strategy DeCK to enhance ICE in overcoming stubborn knowledge. Figure 4 illustrates the process of using DeCK to handle the editing case shown in Figure 1 (right). DeCK can be formalized as follows. Using $\text{I\kern-1.49994ptP}(x_{t})$ to represent $\text{I\kern-1.49994ptP}(x_{t}|x_{<t})$ for notational brevity, we compute the probability of the next token by,
@@
Figure 4: Illustration of DeCK enhancing ICE to overcome stubborn knowledge. During decoding, DeCK contrasts the enhanced ICE distribution with the original distribution to highlight new knowledge, inducing LLMs to generate edited facts during the process of reasoning about the answer to input questions using chain-of-though (CoT).
@@ 5.4 Ablation Study
We conduct ablation experiments on the key components of our DeCK. Table 6 shows how the contrasting coefficient introduced in Equation 6 affects DeCK’s performance. DeCK is highly sensitive to the contrasting coefficient. If $\gamma$ is too large, it can excessively amplify unreasonable token probabilities, significantly reducing DeCK’s performance, even below that of the original ICE. Table 6 demonstrates that the editing signal enhancement introduced in Section 4.1 can consistently enhance DeCK’s performance. This is because it ensures that the enhanced edited knowledge is not filtered out by 7.
Scale $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.5$

### v2

@@ Abstract
The knowledge within large language models (LLMs) may become outdated quickly. While in-context editing (ICE) is currently the most effective method for knowledge editing (KE), it is constrained by the black-box modeling of LLMs and thus lacks interpretability. Our work aims to elucidate the superior performance of ICE in KE by analyzing the impacts of in-context new knowledge on token-wise distributions. We observe that despite a significant boost in logits of the new knowledge, the performance of ICE is still hindered by stubborn knowledge. Stubborn knowledge refers to facts that have gained excessive confidence during pretraining, making them hard to edit effectively. To address this issue and further enhance the performance of ICE, we propose a novel approach termed Decoding by Contrasting Knowledge (DeCK). DeCK derives the distribution of the next token by contrasting the logits obtained from the newly edited knowledge guided by ICE with those from the unedited parametric knowledge. Our experiments consistently demonstrate that DeCK enhances the confidence of LLMs in edited facts. For instance, it improves the performance of LLaMA3-8B-instruct on MQuAKE by up to 219%, demonstrating its capability to strengthen ICE in the editing of stubborn knowledge. DeCK can be easily integrated into any ICE method as a decoding component to enhance editing capabilities. Our work paves the way to develop both effective and accountable KE methods for LLMs.  [^1] The source code is available at: https://github.com/byronBBL/DeCK.
@@ 1 Introduction
Unlike the high computational resource burden incurred by retraining from scratch, knowledge editing (KE)  (Sinitsin et al.,, 2020; De Cao et al.,, 2021; Zhu et al.,, 2020; Mitchell et al.,, 2022; Yao et al.,, 2023) has been proposed as an efficient means to update the knowledge of LLMs. They aim to edit knowledge by incrementally injecting or modifying facts.
As LLMs demonstrate increasingly powerful in-context learning capabilities, recent research (Madaan et al.,, 2022; Zhong et al.,, 2023; Zheng et al.,, 2023; Cohen et al.,, 2024; Wang et al.,, 2024; Mei et al.,, 2024) has delved into easier and efficient methods for in-context editing (ICE), aiming to directly guide frozen LLMs in generating text with new knowledge through contextual prompts. Figure 1 (left) illustrates an example of successful editing using ICE. These ICE methods showcasing state-of-the-art performance without the need to alter internal model parameters, indicate the promising potential of modeling LLMs as black boxes for ICE guided by external contexts.
@@
However, as illustrated in Figure 1 (middle), there still exist deeply entrenched pieces of knowledge in LLMs that are difficult for ICE to modify, which we refer to as stubborn knowledge. We argue that LLMs, through extensive pre-training, have developed strong confidence in certain facts, making them difficult to alter solely through external contextual prompts (Bi et al.,, 2024). Therefore, despite the fact that the sophisticated methods such as enhancing retrieval (Shi et al.,, 2024), checking conflict (Zhong et al.,, 2023), and guiding reasoning (Wang et al.,, 2024) can enhance the overall performance of ICE, relying on these external methods still cannot genuinely improve the foundational capability for editing individual stubborn knowledge.
In this work, we focus on enhancing the state-of-the-art KE method, ICE, to reduce the negative impacts from the stubborn knowledge in LLMs. First, we observe the impact of the in-context new knowledge in ICE on LLMs from the perspective of LLMs’ token-level distributions. We find that incorporating this new knowledge significantly increases the predicted probability of generating edited facts during the decoding process. A deeper exploration of the failed cases reveals the reasons why stubborn knowledge is difficult to edit. Despite the significant improvement in the logits of new knowledge achieved by ICE, there persists a small gap between new knowledge and parametric knowledge, where parametric knowledge refers to the original unedited knowledge in LLMs.
Building upon the insights gained from above observations, we introduce a new decoding technique called Decoding by Contrasting Knowledge (DeCK) to enhance LLMs’ confidence in edited facts for better editing of stubborn knowledge. DeCK consists of two components: (1) an editing enhancement module that improves attention to new knowledge, thus preventing it from being filtered out during contrastive decoding, and (2) a contrastive decoding strategy that compares the logical distributions after in-context editing with the original parametric logical distributions to predict the next token.
Overall, our contributions can be summarized by three points. First, as far as we know, we are the first to elucidate superior performance of ICE on the KE from a model interpretability perspective. Second, we find that stubborn knowledge significantly impacts the performance of ICE, and we propose DeCK to boost confidence in editing facts, enhancing ICE to overcome it. Third, extensive experiments on MQuAKE indicate that our DeCK can effectively enhance the performance of ICE without altering the internal model or modifying external prompts. DeCK can be easily integrated into any ICE method as a decoding component to enhance editing capabilities. Our work paves the way to develop the both effective and accountable KE methods for LLMs.
@@ 3 Deep Insights into ICE through Decoding Perspectives
With $\phi(\cdot)$ replacing the affine layer to predict the probability of the next token over the vocabulary set $\mathcal{V}$, we can obtain a simplified representation of Equation (1). Given a sequence of tokens $\mathcal{X}_{E}=\{x_{1}^{(E)},x_{2}^{(E)},...,x_{m-1}^{(E)}\}$, which includes guidance from an editing prompt, such as "Assume Tesla’s CEO is Jack Ma", we compute the probability of next token $x^{(E)}_{m}$ with editing guidance as follows:
@@
We can also represent the parametric probability distribution $\text{I\kern-1.49994ptP}^{B}(x^{(B)}_{n}|x^{(B)}_{<n})$ by considering only the token sequence $\mathcal{X}_{B}$ containing the original question prompt without any editing content. The distribution $\text{I\kern-1.49994ptP}^{E}(x^{(E)}_{m}|x^{(E)}_{<m})$ also reflects the feedback from the introduction of external knowledge, while $\text{I\kern-1.49994ptP}^{B}(x^{(B)}_{n}|x^{(B)}_{<n})$ solely represents the response of LLMs based on their parametric knowledge to the question.
@@ 3.1 How ICE Effectively Edits Knowledge in LLMs?
Although the ICE methods on LLMs (Zhong et al.,, 2023; Cohen et al.,, 2024; Wang et al.,, 2024) have demonstrated promising performance, they all rely on the black-box modeling of LLMs for editing, and the internal mechanisms behind their effectiveness remain unclear. In this subsection, we delve into the intrinsic reasons behind the superior performance of ICE on KE.
We designed dedicated experiments to capture the logits output of knowledge that would be influenced by the edit. A striking observation in Figure 2 is that the introduction of new knowledge through ICE leads to a significant rightward shift in the probability distribution of the new knowledge, while the logits for parametric knowledge remain largely unchanged or decrease to some extent. This suggests that ICE significantly enhances the logits of new knowledge while having minimal impact on parametric knowledge. Additionally, the number of top-ranked positions for new knowledge significantly increases after ICE, with the majority surpassing that of parametric knowledge. This indicates that the in-context new knowledge can improve the confidence of LLMs in editing facts, thereby prompting responses with the edited answers.
@@ 3.2 Inherent Challenges of Stubborn Knowledge
While ICE has significantly boosted the confidence of LLMs in new knowledge, we find that there are still instances where certain new knowledge ranks prominently but not as the top-1, as illustrated in Figure 2. We term this phenomenon "stubborn knowledge", which refers to cases where editing fails due to either an excessive confidence in existing parametric knowledge or insufficient confidence in new knowledge. The edit cases in Figure 3 deeply reveals the failed pattern for ICE in addressing stubborn knowledge, which happens when there is still an extremely small gap compared to the parametric knowledge after editing, despite the significant increase in new knowledge logits induced by the editing prompt. Taking the last case as an example, after editing, the new knowledge "English" lags behind the parametric knowledge "French" by only 0.516 in terms of logical distribution, illustrating how a minor gap leads to editing failure. This indicates the intrinsic reasons for the failure of black-box ICE methods to edit stubborn knowledge in LLMs in most cases.
@@ 4 DeCK: Enhancing LLMs’ Confidence on Edited Facts
Inspired by the observations in Section 3, we design our novel decoding strategy DeCK to enhance ICE in overcoming stubborn knowledge. Figure 4 illustrates the process of using DeCK to handle the stubborn knowledge case shown in Figure 1 (right). DeCK can be formalized as follows. Using $\text{I\kern-1.49994ptP}(x_{t})$ to represent $\text{I\kern-1.49994ptP}(x_{t}|x_{<t})$ for notational brevity, we compute the probability of the next token by,
@@
Figure 4: Illustration of DeCK enhancing ICE to edit the stubborn knowledge. During decoding, DeCK contrasts the enhanced ICE distribution with the original distribution to highlight new knowledge, inducing LLMs to generate edited facts using chain-of-though (CoT) (Wei et al.,, 2022) during the reasoning process for answering input questions.
@@ 5.4 Ablation Study
We conduct ablation experiments on the key components of our DeCK. Table 6 shows how the contrasting coefficient introduced in Equation 6 affects DeCK’s performance. DeCK is highly sensitive to the contrasting coefficient. If $\gamma$ is too large, it can excessively amplify unreasonable token probabilities, significantly reducing DeCK’s performance, even below that of the original ICE. Table 6 demonstrates that the editing signal enhancement introduced in Section 4.1 can consistently enhance DeCK’s performance. This is because it ensures that the enhanced edited knowledge is not filtered out by Equation 7.
Size $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.5$